# Iron Sharpens Iron: Defending Against Attacks in Machine-Generated Text Detection with Adversarial Training

### v1

@@ Abstract
Machine-generated Text (MGT) detection is crucial for regulating and attributing online texts. While the existing MGT detectors achieve strong performance, they remain vulnerable to simple perturbations and adversarial attacks. To build an effective defense against malicious perturbations, we view MGT detection from a threat modeling perspective, that is, analyzing the model’s vulnerability from an adversary’s point of view and exploring effective mitigations. To this end, we introduce an adversarial framework for training a robust MGT detector, named GREedy Adversary PromoTed DefendER (Greater). The Greater consists of two key components: an adversary Greater-A and a detector Greater-D. The Greater-D learns to defense against the adversarial attack from Greater-A and generalizes the defense to other attacks. Greater-A identifies and perturbs the critical tokens in embedding space, along with greedy search and pruning to generate stealthy and disruptive adversarial examples. Besides, we update the Greater-A and Greater-D synchronously, encouraging the Greater-D to generalize its defense to different attacks and varying attack intensities. Our experimental results across 9 text perturbation strategies and 5 adversarial attacks show that our Greater-D reduces the Attack Success Rate (ASR) by 10.61% compared with SOTA defense methods while our Greater-A is demonstrated to be more effective and efficient than SOTA attack approaches.
@@ 1 Introduction
Figure 1: Performance drop of different MGT detectors and defense methods under text perturbation [^2] The detectors include DetectGPT Mitchell et al. (2023), LogRank Su et al. (2023), fine-tuned xlm-roberta-base Liu et al. (2019), fine-tuned Albert-large Lan et al. (2019), and Unbiased Watermark Hu et al. (2023). The defense methods contain Text-RS Zhang et al. (2024b), CERT-ED Huang et al. (2024), VAT Miyato et al. (2016), TAVAT Li et al. (2021), and GREATER-D (ours)..
@@
Motivation. We rely on threat modeling to advance the robustness of the MGT detectors against perturbation and adversarial attacks. As the proverb says ’Iron sharpens iron’, we focus on constructing powerful adversarial examples which mislead the prediction of the detector to facilitate the post-training of MGT detectors and defend against different attacks. Existing text perturbation strategies Wang et al. (2024a) adjust token distribution without accessing information from the target MGT detector, resulting in low-quality and non-targeted adversarial examples. The adversarial attacks are only effective in white-box setting Yoo and Qi (2021) or require excessive queries to target detectors Hu et al. (2024); Yu et al. (2024); Liu et al. (2024a). Moreover, Wang et al. (2024b) find that the defense built by adversarial training cannot generalize well to the attacks on which it was not originally trained. To overcome these limitations, we propose an efficient adversarial training framework that works in a black-box setting and builds generalizable defense against a wide variety of perturbations and attacks for MGT detectors.
Our Work. In this paper, we propose an adversarial framework for training robust MGT detector, namely GREedy Adversary PromoTed DefendeR. (Greater). Greater consists of an adversary (Greater-A) and a detector (Greater-D). The Greater-D learns to discern MGTs from the human-written texts (HWTs), while the Greater-A, which queries the output of the detector, aims to imply minimum perturbation on MGTs to deceive the detector. Restricted by the scenario where only outputs from the target detector are available, we use an open-sourced surrogate model to retrieve gradient information to identify important tokens in the prediction. Afterwards, we introduce a gradient ascent perturbation on the embedding of MGTs from the surrogate model to enhance both the quality and stealthiness of generated adversarial text. To reduce the number of queries needed for building effective adversarial examples, we design a greedy search and pruning strategy. In the training stage, we update the Greater-A and Greater-D in the same training step so that Greater-D learns from a curriculum of adversarial examples to generalize its defense. The experiment results demonstrate that our method achieves an average ASR of 5.75% against various attacks, which is 10.61% lower compared to the SOTA defense methods. We also find that Greater-A achieves the most effective attack, achieving an ASR of 96.58%, which surpasses SOTA attack methods by 8.45% while requiring 4 times fewer queries.
@@
  • Outstanding Performance. Testing results across 14 attack methods demonstrate that our detector outperforms 8 existing SOTA defense methods in robustness, while our adversary achieves significant improvements in both attack efficiency and effectiveness compared to 13 SOTA attack approaches.
@@ 2 Related Work
Adversarial Training. Adversarial training aims to optimize the model toward maintaining correct predictions to adversarial examples that are misleading data constructed for malicious purposes. Earlier works first augment the training set with adversarial examples for defense against specific attacks Huang et al. (2024); Zeng et al. (2023). These methods are shown to be hard to generalize to unseen attacks Wang et al. (2024b). Yoo and Qi (2021); Li et al. (2021) update the adversary and the target model in the same step to generalize the defense to unseen attacks. However, they rely on the availability of explicit first-order gradient, which is not applicable in the real-world case. Different from previous works, we propose an effective adversarial training framework for MGT detectors that builds a generalizable defense against a variety of attacks in the black-box setting.
@@ 4.1.2 Replace & Refine
To further validate the efficiency of our method, we provide a detailed theoretical analysis of query complexity and perturbation rate in Greater-A in Appendix D.
@@ 4.3 Adversarial Training
Training Process. We update the adversary and the detector alternatively in the same training step. Specifically, at each training step $t$, we first update the adversary with loss function (10) while keeping the detector frozen. The updated adversary generates adversarial example $\tilde{X}_{i}$ in the current step. Then we update the detector with loss function (12). We detail the adversarial training process in form of pseudocode in Appendix E.
@@ 5 Experiment Results
F.t.XLM-RoBERTa-Base EP PP CERT-ED RanMask Text-RS Text-CRS VAT TAVAT Greater-D
Text Perturbation Mixed Edit ASR(%$\downarrow$ 34.65 4.58 32.33 16.74 17.00 22.81 15.34 33.19 37.50 8.85
Paraphrasing ASR(%$\downarrow$ 70.58 54.65 6.27 32.10 40.20 59.58 26.67 67.98 65.90 3.45
Code-switching MF ASR(%$\downarrow$ 50.58 38.37 34.08 29.19 27.78 48.80 27.15 36.71 47.52 1.13
Code-switching MR ASR(%$\downarrow$ 47.91 31.23 30.21 5.30 10.80 16.98 14.36 38.03 46.46 1.02
Human Obfuscation ASR(%$\downarrow$ 18.42 22.19 27.00 13.64 18.86 18.05 15.24 25.35 24.17 0.86
Emoji-cogen ASR(%$\downarrow$ 32.19 46.55 44.17 11.41 22.23 17.72 27.10 52.35 40.33 0.47
Typo-cogen ASR(%$\downarrow$ 60.10 61.57 59.72 27.29 38.82 37.09 44.79 70.26 63.04 1.08
ICL ASR(%$\downarrow$ 1.40 1.41 1.30 1.72 0.83 1.89 0.67 1.13 1.88 0.20
Prompt Paraphrasing ASR(%$\downarrow$ 0.00 0.69 0.00 0.70 0.00 0.00 0.00 0.00 0.00 0.23
CSGen ASR(%$\downarrow$ 25.44 22.81 6.14 10.65 1.70 2.41 23.95 26.88 27.52 0.00
@@
Adversarial Attack PWWS ASR(%$\downarrow$ 61.45 48.47 49.06 11.07 14.83 16.13 19.15 53.56 62.68 5.91
@@
GREATER-A(ours) ASR(%$\downarrow$ 96.58 87.08 84.34 62.17 63.58 75.25 82.29 89.26 85.02 46.08
@@
Total Avg. ASR(%$\downarrow$ 45.90 39.81 37.22 16.36 19.85 24.21 23.40 44.17 49.63 5.75
Table 1: Performance of defense methods under different attacks. The best results are highlighted in green background. {*} means that Code-switching MF and Code-switching MR are two variations of Code-switching method.
@@ 5.1 Defense Performance for Greater-D
Experiment Setting. We evaluate our defense model Greater-D against 14 text perturbation and adversarial attack methods, whose detailed introductions are outlined in the Appendix A.5. The competitors include i) data augmentation methods: Editing Pretrained (EP) Wang et al. (2024b), Paraphrasing Pretrained (PP) Wang et al. (2024b), CERT-ED Huang et al. (2024), RanMask Zeng et al. (2023), Text-RS Zhang et al. (2024b), Text-CRS Zhang et al. (2024a). ii) adversarial training methods: Virtual Adversarial Training (VAT) Miyato et al. (2016), Token Aware Virtual Adversarial Training (TAVAT) Li et al. (2021). Detailed introduction and implementation are presented in Appendix A.4.1 and A.2. The dataset we use is presented in Appendix A.1.
Experiment results. We present the defense performance of Greater-D in Table 1 and unveil the following three key insights: 1) Best defense performance. Our method exhibits the best defense performance against different attacks among all competitors. The average Attack Success Rate (ASR) drops to 1.73% for text perturbation attack and 13.78% for adversarial attack, respectively, which is lower by 13.15% and 5.54% compared with the second-best defense method. 2) Most effort needed for adversarial attack. We observe that it takes adversarial attack more resources to conduct a successful attack to Greater-D. 843.29 queries are required in average, which is 31.36 more than other defense methods. Moreover, the average ASR against Greater-D is only 13.78% for adversarial attacks. It illustrates that Greater-D makes adversarial attacks both inefficient and ineffective. 3) Generalized Defense to Different Attacks. We notice that Greater-D significantly reduces the ASR of different kinds of attacks even though it is trained with Greater-A. As an exemplary method, EP performs better than Greater-D when defending Mixed Edit Attack but cannot withstand other attacks. The defense against a wide variety of attacks demonstrates the generalized defense effect of Greater-D.
@@ 7 Conclusion
In this paper, we proposed an adversarial training framework GREedy Adversary PromoTed DefendER (Greater) to enhance the robustness of MGT detector under different text perturbation and adversarial attacks. We design a novel attack strategy for the adversary including Identify & Perturb and Replace & Refine to construct effective adversarial examples efficiently. In Greater, we update the adversary and the detector alternatively in the same training step for better defense generalization. Our experiment results demonstrate the efficacy of our detector Greater-D under 14 attacks along with the leading performance of adversary Greater-A compared with 13 methods. The discussion on the relationship between attack strength and defense performance reveals the importance of a powerful adversary in adversarial training for robust MGT detectors.
@@
Despite the promising results achieved by Greater, there are two primary limitations to this study. First, our defense method can generalize to different attacks but its application on texts in different languages remains a challenge. Second, the computational cost of training the adversarial framework, particularly for the adversary and detector, is substantial, requiring significant hardware resources that could limit its deployment in resource-constrained settings. Third, the detector presented in our work is only able to attribute the origin of the text on the document-level. However, more works are needed for fine-grained (e.g., sentence-level, token-level) detection in Human-AI co-authored texts.

### v2

@@ Abstract
Machine-generated Text (MGT) detection is crucial for regulating and attributing online texts. While the existing MGT detectors achieve strong performance, they remain vulnerable to simple perturbations and adversarial attacks. To build an effective defense against malicious perturbations, we view MGT detection from a threat modeling perspective, that is, analyzing the model’s vulnerability from an adversary’s point of view and exploring effective mitigations. To this end, we introduce an adversarial framework for training a robust MGT detector, named GREedy Adversary PromoTed DefendER (Greater). The Greater consists of two key components: an adversary Greater-A and a detector Greater-D. The Greater-D learns to defend against the adversarial attack from Greater-A and generalizes the defense to other attacks. Greater-A identifies and perturbs the critical tokens in embedding space, along with greedy search and pruning to generate stealthy and disruptive adversarial examples. Besides, we update the Greater-A and Greater-D synchronously, encouraging the Greater-D to generalize its defense to different attacks and varying attack intensities. Our experimental results across 10 text perturbation strategies and 6 adversarial attacks show that our Greater-D reduces the Attack Success Rate (ASR) by 0.67% compared with SOTA defense methods while our Greater-A is demonstrated to be more effective and efficient than SOTA attack approaches. Codes and dataset are available in https://github.com/Liyuuuu111/GREATER.
@@ 1 Introduction
Figure 1: Performance drop of different MGT detectors and defense methods under text perturbation [^2] The detectors include DetectGPT Mitchell et al. (2023), LogRank Su et al. (2023), fine-tuned xlm-roberta-base Liu et al. (2019), fine-tuned Albert-large Lan et al. (2019), and Unbiased Watermark Hu et al. (2023b). The defense methods contain Text-RS Zhang et al. (2024b), CERT-ED Huang et al. (2024), VAT Miyato et al. (2016), TAVAT Li et al. (2021), and GREATER-D (ours)..
@@
Motivation. We rely on threat modeling to advance the robustness of the MGT detectors against perturbation and adversarial attacks. As the proverb says ’Iron sharpens iron’, we focus on constructing powerful adversarial examples which mislead the prediction of the detector to facilitate the post-training of MGT detectors and defend against different attacks. Existing text perturbation strategies Wang et al. (2024a) adjust token distribution without accessing information from the target MGT detector, resulting in low-quality and non-targeted adversarial examples. The adversarial attacks are only effective in white-box setting Yoo and Qi (2021) or require excessive queries to target detectors Hu et al. (2024); Yu et al. (2024); Liu et al. (2024a). Moreover, Wang et al. (2024c) find that the defense built by adversarial training cannot generalize well to the attacks on which it was not originally trained. To overcome these limitations, we propose an efficient adversarial training framework that works in a black-box setting and builds generalizable defense against a wide variety of perturbations and attacks for MGT detectors.
Our Work. In this paper, we propose an adversarial framework for training robust MGT detector, namely GREedy Adversary PromoTed DefendeR. (Greater). Greater consists of an adversary (Greater-A) and a detector (Greater-D). The Greater-D learns to discern MGTs from the human-written texts (HWTs), while the Greater-A, which queries the output of the detector, aims to imply minimum perturbation on MGTs to deceive the detector. Restricted by the scenario where only outputs from the target detector are available, we use an open-sourced surrogate model to retrieve gradient information to identify important tokens in the prediction. Afterwards, we introduce a gradient ascent perturbation on the embedding of MGTs from the surrogate model to enhance both the quality and stealthiness of generated adversarial text. To reduce the number of queries needed for building effective adversarial examples, we design a greedy search and pruning strategy. In the training stage, we update the Greater-A and Greater-D in the same training step so that Greater-D learns from a curriculum of adversarial examples to generalize its defense. The experiment results demonstrate that our method achieves an average ASR of 5.53% against various attacks, which is 0.67% lower compared to the SOTA defense methods. We also find that Greater-A achieves the most effective attack, achieving an ASR of 96.58%, which surpasses SOTA attack methods by 8.45% while requiring 4 times fewer queries.
@@
  • Outstanding Performance. Testing results across 16 attack methods demonstrate that our detector outperforms 10 existing SOTA defense methods in robustness, while our adversary achieves significant improvements in both attack efficiency and effectiveness compared to 13 SOTA attack approaches.
@@ 2 Related Work
Adversarial Training. Adversarial training aims to optimize the model toward maintaining correct predictions against adversarial examples that are misleading data constructed for malicious purposes. Earlier works first augment the training set with adversarial examples for defense against specific attacks Huang et al. (2024); Zeng et al. (2023). These methods are shown to be hard to generalize to unseen attacks Wang et al. (2024c). Yoo and Qi (2021) update the adversary and the target model in the same step to generalize the defense to unseen attacks. However, they rely on the availability of explicit first-order gradients, which is not applicable in the real-world case. OUTFOX Koike et al. (2024) identifies adversarial attack with in-context learning, which requires the demonstrations of adversarial samples as prompts. RADAR Hu et al. (2023a) uses a paraphraser as the adversary. Both OUTFOX and RADAR are designed to defend against known attacks. Different from previous works, we propose an effective adversarial training framework for MGT detectors that builds a generalizable defense against a variety of attacks in the black-box setting.
@@ 4.1.2 Replace & Refine
To further validate the efficiency of our method, we provide a detailed theoretical analysis of query complexity and perturbation rate in Greater-A in Appendix E.
@@ 4.3 Adversarial Training
Training Process. We update the adversary and the detector alternatively in the same training step. Specifically, at each training step $t$, we first update the adversary with loss function (10) while keeping the detector frozen. The updated adversary generates adversarial example $\tilde{X}_{i}$ in the current step. Then we update the detector with loss function (12). We detail the adversarial training process in form of pseudocode in Appendix F.
@@ 5 Experiment Results
F.t.XLM-RoBERTa-Base EP PP CERT-ED RanMask Text-RS Text-CRS VAT TAVAT RADAR OUTFOX GREATER-D
Text Perturbation Mixed Edit ASR(%$\downarrow$ 34.65 4.58 32.33 16.74 17.00 22.81 15.34 33.19 37.50 18.76 11.33 8.85
HMGC ASR(%$\downarrow$ 30.00 6.53 27.34 12.52 11.74 16.58 13.37 25.19 31.50 8.53 3.74 2.28
Paraphrasing ASR(%$\downarrow$ 70.58 54.65 6.27 32.10 40.20 59.58 26.67 67.98 65.90 2.08 14.96 3.45
Code-switching MF ASR(%$\downarrow$ 50.58 38.37 34.08 29.19 27.78 48.80 27.15 36.71 47.52 32.57 5.46 1.13
Code-switching MR ASR(%$\downarrow$ 47.91 31.23 30.21 5.30 10.80 16.98 14.36 38.03 46.46 14.14 5.37 1.02
Human Obfuscation ASR(%$\downarrow$ 18.42 22.19 27.00 13.64 18.86 18.05 15.24 25.35 24.17 16.26 5.74 0.86
Emoji-cogen ASR(%$\downarrow$ 32.19 46.55 44.17 11.41 22.23 17.72 27.10 52.35 40.33 33.13 7.72 0.47
Typo-cogen ASR(%$\downarrow$ 60.10 61.57 59.72 27.29 38.82 37.09 44.79 70.26 63.04 41.52 9.29 1.08
ICL ASR(%$\downarrow$ 1.40 1.41 1.30 1.72 0.83 1.89 0.67 1.13 1.88 2.77 0.59 0.20
Prompt Paraphrasing ASR(%$\downarrow$ 0.00 0.69 0.00 0.70 0.00 0.00 0.00 0.00 0.00 0.94 0.65 0.23
CSGen ASR(%$\downarrow$ 25.44 22.81 6.14 10.65 1.70 2.41 23.95 26.88 27.52 17.07 3.40 0.00
@@
Adversarial Attack PWWS ASR(%$\downarrow$ 61.45 48.47 49.06 11.07 14.83 16.13 19.15 53.56 62.68 12.58 - 5.91
@@
T-PGD ASR(%$\downarrow$ 49.90 34.15 25.28 8.52 12.07 13.15 14.92 42.14 45.58 7.79 - 5.61
[…]
GREATER-A(ours) ASR(%$\downarrow$ 96.58 87.08 84.34 62.17 63.58 75.25 82.29 89.26 85.02 71.04 - 46.08
@@
Total Avg. ASR(%$\downarrow$ 45.20 37.52 35.93 15.67 18.91 23.11 22.31 42.93 48.33 18.55 6.20 5.53
Table 1: Performance of defense methods under different attacks. The best results are highlighted in green background. {*} means that Code-switching MF and Code-switching MR are two variations of Code-switching method. Since the detector in OUTFOX is a closed-source LLM, it cannot be attacked by adversarial attack. Therefore, we only evaluate OUTFOX under text perturbation attacks.
@@ 5.1 Defense Performance for Greater-D
Experiment Setting. We evaluate our defense model Greater-D against 16 text perturbation and adversarial attack methods, whose detailed introductions are outlined in the Appendix A.5. The competitors include i) data augmentation methods: Editing Pretrained (EP) Wang et al. (2024c), Paraphrasing Pretrained (PP) Wang et al. (2024c), CERT-ED Huang et al. (2024), RanMask Zeng et al. (2023), Text-RS Zhang et al. (2024b), Text-CRS Zhang et al. (2024a). ii) adversarial training methods: Virtual Adversarial Training (VAT) Miyato et al. (2016), Token Aware Virtual Adversarial Training (TAVAT) Li et al. (2021), RADAR Hu et al. (2023a), OUTFOX Koike et al. (2024). Detailed introduction and implementation are presented in Appendix A.4.1 and A.2. The dataset we use is presented in Appendix A.1.
Experiment results. We present the defense performance of Greater-D in Table 1 and unveil the following three key insights: 1) Best defense performance. Our method exhibits the best defense performance against different attacks among all competitors. The average Attack Success Rate (ASR) drops to 1.78% for text perturbation attacks and 12.42% for adversarial attacks, which are lower than the second-best methods, RADAR (6.20%) and CERT-ED (17.52%), by 4.42% and 5.10%, respectively. 2) Most effort needed for adversarial attack. We observe that it takes adversarial attack more resources to conduct a successful attack to Greater-D. 840.17 queries are required on average, which is 30.29 more than other defense methods. Moreover, the average ASR against Greater-D is only 12.42% for adversarial attacks. It illustrates that Greater-D makes adversarial attacks both inefficient and ineffective. 3) Generalized Defense to Different Attacks. We notice that Greater-D significantly reduces the ASR of different kinds of attacks even though it is trained with Greater-A. As an exemplary method, EP performs better than Greater-D when defending Mixed Edit Attack but cannot withstand other attacks. The defense against a wide variety of attacks demonstrates the generalized defense effect of Greater-D.
@@ 6.1 Impact of Attack Strength of Greater-A on Greater-D
6.2 Defense Adaptation to Different Backbones
Model Metric Baseline CERT-ED Greater-D
ALBERT Base (12M) ASR(%$\downarrow$ 63.58 46.71 35.62
ALBERT Large (18M) ASR(%$\downarrow$ 97.14 45.45 43.40
DeBERTa Base (86M) ASR(%$\downarrow$ 22.95 29.37 4.10
DeBERTa Large (304M) ASR(%$\downarrow$ 51.72 25.07 10.25
RoBERTa Base (125M) ASR(%$\downarrow$ 81.88 26.78 15.32
RoBERTa Large (355M) ASR(%$\downarrow$ 30.21 34.66 5.38
XLM-RoBERTa Large (561M) ASR(%$\downarrow$ 81.99 40.04 1.73
6.3 Impact of Surrogate Model of Greater-A
We use different surrogate model and dataset for training GREATER-A to demonstrate that GREATER-A’s performance is independent of surrogate model selection and training data. Specifically, we use RoBERTa-Large and GPT2 as surrogate models, and use SemEval (same dataset for training GREATER-D, abbreviated as IND) and M4 (Wang et al., 2024b) (different dataset from training GREATER-D, abbreviated as IND) to validate our claim. The results are shown in Table 4.
Surrogate Model Dataset (Type) Avg Queries ASR (%) Pert. (%) $\Delta$PPL ↓ USE $\Delta r$
RoBERTa-Large SemEval (IND) 62.63 96.58 7.26 35.22 0.9506 9.21
RoBERTa-Large M4 (OOD) 65.77 95.72 7.33 38.84 0.9503 9.65
GPT2 SemEval (IND) 65.60 96.17 7.29 36.87 0.9550 12.73
GPT2 M4 (OOD) 63.80 96.36 7.27 37.14 0.9536 10.64
Table 4: The attack results of different surrogate models and datasets that the surrogate models are trained with. The best result in each group is highlighted with a green background.
As shown in Table 4, the attack effectiveness remains stable regardless of the surrogate model or training dataset. This consistency in attack success rates indicates that GREATER-A’s performance is not sensitive to changes in the surrogate model or training data, supporting the claim that the use of a surrogate model from the same family as the target model does not leak crucial information.
@@ 7 Conclusion
In this paper, we proposed an adversarial training framework GREedy Adversary PromoTed DefendER (Greater) to enhance the robustness of MGT detector under different text perturbation and adversarial attacks. We design a novel attack strategy for the adversary including Identify & Perturb and Replace & Refine to construct effective adversarial examples efficiently. In Greater, we update the adversary and the detector alternatively in the same training step for better defense generalization. Our experiment results demonstrate the efficacy of our detector Greater-D under 16 attacks along with the leading performance of adversary Greater-A compared with 13 methods. The discussion on the relationship between attack strength and defense performance reveals the importance of a powerful adversary in adversarial training for robust MGT detectors.
@@
Despite the promising results achieved by Greater, there are three primary limitations to this study. First, our defense method can generalize to different attacks but its application on texts in different languages remains a challenge. Second, the computational cost of training the adversarial framework, particularly for the adversary and detector, is substantial, requiring significant hardware resources that could limit its deployment in resource-constrained settings. Third, the detector presented in our work is only able to attribute the origin of the text on the document-level. However, more works are needed for fine-grained (e.g., sentence-level, token-level) detection in Human-AI co-authored texts.
@@
Acknowledgment
We thank all the reviewers and the area chair for their helpful feedback, which aided us in greatly improving the paper. This work is supported by National Key R&D Program (2023YFB3107400), National Natural Science Foundation of China (62272371, 62103323, U24B20185, T2442014, 62161160337, 62132011, U21B2018), Initiative Postdocs Supporting Program (BX20190275, BX20200270), China Postdoctoral Science Foundation (2019M663723, 2021M692565), Fundamental Research Funds for the Central Universities under grant (xzy012024144), and Shaanxi Province Key Industry Innovation Program (2023-ZDLGY-38, 2021ZDLGY01-02). Thanks to the New Cornerstone Science Foundation and the Xplorer Prize.